# Jailbreak Attack Initializations as Extractors of Compliance Directions

### v2

@@ 4.2 Experimental Results
In Figure 3, we present the correlation between $LFS$ and the number of optimization steps to success on the $GCG$ attack over the Llama-2 model. Our results entail a clear positive correlation, which supports our assumption of $LFS$ indicating the initializations’ effectiveness. Moreover, nearly all $25\text{-}CRI$ initializations present a low $LFS$ and require relatively few steps to reach success, presenting the initialization set’s effectiveness and the enhancement of corresponding attacks. In contrast, other initialization methods require substantially more computational resources to succeed, and there is no clear indication of the benefit of the baseline initialization over the uninformed initializations. These indications are supported by a similar experiment over the Vicuna model, which we present in Section C.4.4.
@@
In Figure 4, we present the $ASR$ of the $GCG$ and Embedding attacks depending on the number of steps, over the Llama-2 and Vicuna models. In the Embedding attack, all initializations quickly achieve $100\%$ $ASR$, with the uninformed initializations requiring somewhat more computational resources. This may indicate that these settings are less robust, which aligns with the Embedding attack being considered better performing. In addition, these results emphasize the computational benefit of utilizing informed initializations. Similarly, for $GCG$ on the Vicuna model, all initializations achieve nearly $100\%$ $ASR$, yet with more distinct computational requirements. $CRI$ now requires the fewest optimization steps, with comparable performance for $K=1,5,25$, the baseline approach requires somewhat more, and substantially so for the uninformed approaches. The high $ASR$ again indicates this setting as non-robust, where the computational comparison implies the benefit of utilizing the $LFS$ metric for selecting initializations. In contrast, for $GCG$ on the Llama-2 only $25\text{-}CRI,5\text{-}CRI$ achieve near $100\%$ $ASR$, with all other approaches achieving substantially less. This may indicate that this setting is more robust, where the effective initializations of $25\text{-}CRI,5\text{-}CRI$ aid in circumventing this robustness, and is supported by those two approaches exclusively achieving substantial $ASR$ in the first few steps. The baseline initialization then achieves a higher $ASR$ than $1\text{-}CRI$, which suggests that $1\text{-}CRI$ is inefficient in robust scenarios. This is supported by the corresponding training results of $1\text{-}CRI$, presented in Section C.3, where it achieves negligible $ASR$ over the validation set.
@@
In Table 1, we compare the resulting $ASR$ of the $GCG$ and Embedding attacks over the Llama-2 and Vicuna models. $CRI$ then achieves the best $ASR,MSS,ASS,LFS$ in all the presented settings, substantially out performing other initialization strategies. Similarly to previous indications, $5\text{-}CRI,25\text{-}CRI$ consistently achieves the best or comparable results, and $1\text{-}CRI$ is only effective in non-robust settings. This supports our previous indications that $LFS$-based initialization selection is an effective initialization strategy.
@@ 5 Discussion
This work suggests $CRI$, an initialization framework for gradient-based jailbreak attacks. Our findings suggest that given attacks optimize prompts’ distributions by gradually shifting them along single compliance directions. We then define the $LFS$ metric to indicate progress along this direction and corresponding proximity to the compliance subspace. This metric both guides the optimization of the $CRI$ set and provides our initialization selection criterion. Our results indicate that this approach produces and utilizes effective initializations, which enhance attacks’ $ASR$ and reduce their computational overhead. For example, with the $GCG$ attack on Llama-2, $25\text{-}CRI$ achieves $100\%$ $ASR$ with an $MSS$ of $2$, while randomly selected initializations from the same set only achieves $78\%$ $ASR$ with an $122$ $MSS$. The uninformative initializations then achieve $50\%$ $ASR$ with an $MSS$ of $185$.
@@
The success of our methods supports the assumption that given attacks gradually follow singular directions. This implies that the safeguards of LLMs can be circumvented with relative ease, as only the identification of such directions is required. Moreover, these directions are defined over the input space and could potentially be transferred to other LLMs. Therefore, eliminating these compliance directions is vital to producing robust LLMs. Jailbreak attacks may still persist in models that achieve such. However, they will require distinct optimization for each malicious prompt, significantly increasing their computational overhead.
@@ A.1 Fig-1: Generation of 2D Visualization
The visualization in Figure 5, is a refined variant of Figure 1, is presented under identical experimental conditions. In this figure, each point represents a distinct harmful prompt. The color scheme differentiates prompt types: red denotes clean harmful prompts, orange indicates prompts initialized using $GCG$’s standard method, and cyan corresponds to prompts initialized with our proposed method ($1\text{-}CRI$). To analyze these prompts, we first obtained their embeddings using Llama-2 and determined whether Llama-2 complied with each prompt. Using these embeddings and compliance labels, we trained an SVM classifier. This process yielded a weight vector $w$ and bias $b$, which define the SVM decision function: $\langle w,x\rangle+b$. A negative SVM score indicates a refusal, while a positive score signifies compliance. This score is plotted on the x-axis. The y-axis represents a one-dimensional t-SNE projection. The orange and cyan paths illustrate the optimization trajectories of the attack: the orange path follows $GCG$’s standard initialization, while the cyan path traces the attack starting from our initialization. The x-axis values are computed using the previously trained SVM classifier.
Figure 5: Visualizations of harmful prompts’ initializations and corresponding $GCG$’s attack steps. Disclaimer, $CRI$ is not an attack and "$CRI$ Attack Steps" are for the attack steps when using $CRI$ as an initialization.
@@ B.2 K-CRI algorithm
Algorithm 1 Compliance–Refusal Initialization ($CRI$)
1:  Input: Fine-tuning set $S_{\mathrm{FT}}$, number of clusters $K$, encoder $EN$, universal attack $A^{U}$
@@ B.3 Prompt Clustering
Figure 8: $AdvBench$ (left) and $HarmBench$ (right) prompt clustering for $5\text{-}CRI$, projected to $2D$ using t-SNE.
@@ B.4 CRI Integration Guidelines
To support the broad applicability of the $CRI$ framework, we outline generic integration strategies adaptable to various optimization-based paradigms. These strategies utilize $K\text{-}CRI$ templates to initialize prompts near the compliance subspace, enhancing efficiency while remaining agnostic to specific attack mechanisms. Integration approaches align and adaptable across gradient-based jailbreak attacks such sampling-based, and reinforcement-driven optimization settings. Concrete integration examples are provided in our experiments, in Section 4.
@@
This paradigm employs gradient-based optimization to iteratively refine input prompts, minimizing the negative log probability of eliciting a target response. The $CRI$ framework integrates by initializing prompts utilizing $K$ pre-computed attacks, which serve as starting points for gradient descent, thereby reducing the number of iterations required for convergence. By leveraging the semantic clustering inherent in $CRI$, this approach ensures that initial prompts are optimally positioned within the compliance subspace, enhancing optimization efficiency Zou et al. (2023); Wallace et al. (2019); Shin et al. (2020); Guo et al. (2021); Ebrahimi et al. (2018). Similarly, $CRI$ can also be integrated for extended attack variants of this category, which utilize a scalar energy function Guo et al. (2024) to guide prompt modifications toward model outputs that align with desired behaviors, often without specifying a precise target response. $CRI$ integrates by selecting $K$ attacks that minimize the initial energy score, providing robust initial prompts for untargeted optimization. This strategy capitalizes on $CRI$’s ability to group prompts by semantic similarity, thereby improving generalizability across diverse prompt distributions Liu et al. (2023); Chao et al. (2023); Zhang et al. (2022); Wei et al. (2024); Li et al. (2023a). Results for Zou et al. (2023) can be found in Table 1,
@@
Embedding attacks manipulate prompt representations within the latent space to circumvent refusal mechanisms, using training-time attacks as a baseline to identify vulnerable directions that trigger compliance. $CRI$ integrates by selecting $K$ attacks within the LLMs’ embedding space, followed by selecting the one that minimizes the attack criterion for a given prompt Leo et al. (2024); Qi et al. (2023); Li et al. (2024a); Zhao et al. (2023); Huang et al. (2023), results for Schwinn et al. (2024) appear in Table 1.
@@
These attacks optimize a single transformation applicable across multiple prompts, prioritizing broad generalizability. $CRI$ integrates by training $K$ attacks on clustered subsets of the fine-tuning set, then selecting the optimal template for a given prompt during inference. This approach balances universal applicability with prompt-specific efficacy, leveraging $CRI$’s clustered structure to enhance transformation robustness Zou et al. (2023); Mozes et al. (2021); Liu et al. (2022); Brown et al. (2023); Yang et al. (2024b), results for Zou et al. (2023) can be found in Table 3 for Llama-2 and Table 3 for Vicuna.
@@
These attacks rely on querying the model without gradient access, using output feedback to iteratively refine prompts. $CRI$ integrates by employing $K$ attacks as an initial set, minimizing the number of model interactions required to achieve compliance. The diversity of $CRI$ clusters enhances query efficiency, particularly in resource-constrained black-box settings Zou et al. (2023); Chen et al. (2021); Garg et al. (2022); Li et al. (2023b); Zhang et al. (2024), for results on the $GCG-M$ model transferability Zou et al. (2023) presented in Table 3 for Llama-2-based jailbreak attack and Table 3 for Vicuna-based jailbreak attack.
@@
Beyond standard optimization-based attacks, $CRI$ extends to additional paradigms such as prefix optimization, refusal unlearning, and alignment modulation Jones et al. (2024); Smith et al. (2024); Xu et al. (2023); Lin et al. (2023); Wu et al. (2024a). Its flexible $K$ initialization set framework enables adaptation to diverse objectives while remaining model- and method-agnostic. Moreover, $CRI$ integrates seamlessly with recent optimization improvements, including techniques like BOOST and I-GCG Boost et al. (2024); IGCG et al. (2024), enhancing convergence and success rates by leveraging robust initializations. An example integration of $CRI$ with BOOST, can be found in Figure 25.
@@
Safety-Alignment divides input prompts into compliance and refusal regions—boundaries that adversaries can exploit (Marshall et al., 2024; Baumann, 2024; Huang et al., 2024; Yu et al., 2024b; Chao et al., 2023; Deng et al., 2023; Li et al., 2025; Yan et al., 2024; Xu et al., 2024b; Xiao et al., 2024a). Recent studies suggest refusal direction—a single vector in the residual stream that governs refusal behavior (Arditi et al., 2024). Modifying this direction can toggle the model’s safety responses, prompting offensive uses like CAST’s real-time content control (Lee et al., 2025) or defensive uses like GradSafe’s gradient-based detection (Xie et al., 2024). However, new work represents refusal cones—suggesting multiple interacting refusal directions (Wollschläger et al., 2025).
@@
Early Jailbreak attack approaches utilize uninformative seeds, employing repeated characters or random Gaussian token embeddings (Zou et al., 2023; Hu et al., 2024). Subsequent methods introduced techniques such as Langevin dynamics or beam search (Guo et al., 2024; Sadasivan et al., 2024; Mehrotra et al., 2023). Implicit initialization approaches rely on subtle mechanisms—few-shot distractors, puzzle-like games, or dynamic context shifts—that implicitly guide models toward generating harmful outputs (Xiao et al., 2024b; Chang et al., 2024; Ramesh et al., 2025). Transfer-based initializations leveraging previously successful jailbreak prompts as starting points for subsequent optimizations (Jia et al., 2024; Andriushchenko et al., 2024; Wu et al., 2025b). Explicit initialization methods directly embed crafted injection prompts or obfuscated content (Liu et al., 2024a, 2023; Schwinn et al., 2024; Liu et al., 2024b; Sun et al., 2025; Xu et al., 2025; Jiang et al., 2024; Zhao et al., 2025a). Additional possible enhancements to initialization strategies include employing energy-based objectives (Zhao et al., 2024), prefix-tuning methods (Zhu et al., 2024; Zhou et al., 2024a; Zhang and Wei, 2025; Li et al., 2024d), or gradient-based optimization techniques (Li et al., 2024b; Yu et al., 2024a).
@@ C.1 Experiments Settings
$GCG$ and the Embedding attacks keep the same settings as before.
@@ C.1.3 Attack Settings
For the $AutoDAN$ attack variants ($GA$, $HGA$) Liu et al. (2023), we utilize the same reference file filled with hand-crafted jailbreaks, and perform $100$ optimization steps.

### v3

@@ 4.2 Experimental Results
In Figure˜4, we present the distribution of per-prompt Spearman correlations between the $LFS$ metric and the number of steps to a successful attack, on Llama-2. For each prompt, we utilize 18 different initializations to calculate both measures, including the standard, random, baseline, and the initializations used in a separate ablation study over the number of training steps as in Figures˜21, 22 and 23. The distribution seems to be heavily skewed to the right, with most per-prompt Spearman correlations clustered at relatively high values (median 0.74), and accompanied by low p-values (median $4.61\times 10^{-4}$). These results confirm a statistically significant, positive monotonic relationship between $LFS$ and steps to success when selecting an initialization for a given prompt, supporting our approach for selecting initializations via the $LFS$ metric. In Figure˜16, the Vicuna distribution also leans to the right, but the median correlation and p-values are somewhat decreased. This may be due to Vicuna being considerably less robust than Llama-2, which aligns with our main results.
@@
In Figure˜5, we present the $ASR$ of the $GCG$ and Embedding attacks depending on the number of steps, over the Llama-2 and Vicuna models. In the Embedding attack, all initializations quickly achieve $100\%$ ASRASR, with the uninformed initializations requiring somewhat more computational resources. This may indicate that these settings are less robust, which aligns with the Embedding attack being considered better performing. In addition, these results emphasize the computational benefit of utilizing informed initializations. Comparable results are presented on newer models in Figures˜11 and 15. Similarly, for $GCG$ on the Vicuna model, all initializations achieve nearly $100\%$ ASRASR, yet with more distinct computational requirements. $CRI$ now requires the fewest optimization steps, with comparable performance for $K=1,5,25$, the baseline approach requires somewhat more, and substantially so for the uninformed approaches. The high $ASR$ again indicates this setting as non-robust, where the computational comparison implies the benefit of utilizing the $LFS$ metric for selecting initializations. In contrast, for $GCG$ on Llama-2, only $25\text{-}CRI,5\text{-}CRI$ achieve near $100\%$ ASRASR, with all other approaches achieving substantially less. This may indicate that this setting is more robust, where the effective initializations of $25\text{-}CRI,5\text{-}CRI$ aid in circumventing this robustness, and is supported by those two approaches exclusively achieving substantial $ASR$ in the first few steps. The baseline initialization then achieves a higher $ASR$ than $1\text{-}CRI$, which suggests that $1\text{-}CRI$ is inefficient in robust scenarios. This is supported by the corresponding training results of $1\text{-}CRI$, presented in Section˜C.3, where it achieves negligible $ASR$ over the validation set.
@@
In Table˜1, we compare the resulting $ASR$ of the $GCG$ and Embedding attacks over the Llama-2 and Vicuna models. $CRI$ then achieves the best $ASR,MSS,ASS,LFS$ in all the presented settings, substantially outperforming other initialization strategies. This is further supported by more models, data, and attacks in Table˜4. Similarly to previous indications, $5\text{-}CRI,25\text{-}CRI$ consistently achieves the best or comparable results, and $1\text{-}CRI$ is only effective in non-robust settings. This supports our previous indications that $LFS$-based initialization selection is an effective initialization strategy.
@@ 5 Discussion
This work suggests $CRI$, an initialization framework for gradient-based jailbreak attacks. Our findings suggest that given attacks optimize prompts’ distributions by gradually shifting them along single compliance directions. We then define the $LFS$ metric to indicate progress along this direction and corresponding proximity to the compliance subspace. This metric both guides the optimization of the $CRI$ set and provides our initialization selection criterion. Our results indicate that this approach produces and utilizes effective initializations, which enhance attacks’ $ASR$ and reduce their computational overhead. For example, with the $GCG$ attack on Llama-2, $25\text{-}CRI$ achieves $100\%$ ASRASR with an $MSS$ of $2$, while randomly selected initializations from the same initialization set only achieve $78\%$ ASRASR with an $122$ MSSMSS, emphasizing the strength of $CRI$’s selection criterion. The uninformative initializations then achieve $50\%$ ASRASR with an $MSS$ of $185$.
@@
The success of our methods supports the assumption that given attacks gradually follow singular directions. This implies that the safeguards of LLMs can be circumvented with relative ease, as only the identification of such directions is required. Moreover, these directions are defined over the input space and could potentially be transferred to other LLMs. Therefore, eliminating these compliance directions is vital to producing robust LLMs. Jailbreak attacks may still persist in models that achieve such. However, they will require distinct optimization for each malicious prompt, significantly increasing their computational overhead. Future studies could leverage $CRI$ as a framework to gauge LLM robustness, specifically by testing how well a model resists attacks with highly effective initializations. This can be measured by evaluating the minimal required computational resources for computing an effective $CRI$ set, i.e., the minimal number of prompts and their optimization overhead required to increase the success of initialized prompts and reduce per-attack computational overhead.
@@ A.1 Fig-1: Generation of 2D Visualization
The visualization in Figure˜6, is a refined variant of Figure˜1, is presented under identical experimental conditions. In this figure, each point represents a distinct harmful prompt. The color scheme differentiates prompt types: red denotes clean harmful prompts, orange indicates prompts initialized using $GCG$’s standard method, and cyan corresponds to prompts initialized with our proposed method ($1\text{-}CRI$). To analyze these prompts, we first obtained their embeddings using Llama-2 and determined whether Llama-2 complied with each prompt. Using these embeddings and compliance labels, we trained an SVM classifier. This process yielded a weight vector $w$ and bias $b$, which define the SVM decision function: $\langle w,x\rangle+b$. A negative SVM score indicates a refusal, while a positive score signifies compliance. This score is plotted on the x-axis. The y-axis represents a one-dimensional t-SNE projection. The orange and cyan paths illustrate the optimization trajectories of the attack: the orange path follows $GCG$’s standard initialization, while the cyan path traces the attack starting from our initialization. The x-axis values are computed using the previously trained SVM classifier.
Figure 6: Visualizations of harmful prompts’ initializations and corresponding GCGGCG’s attack steps. Disclaimer, CRICRI is not an attack and "CRICRI Attack Steps" are for the attack steps when using CRICRI as an initialization.
@@ B.2 K-CRI algorithm
Algorithm 1 Compliance–Refusal Initialization (CRICRI)
1:Input: Fine-tuning set $S_{\mathrm{FT}}$, number of clusters $K$, encoder ENEN, universal attack $A^{U}$
@@ B.3 Prompt Clustering
Figure 9: AdvBenchAdvBench (left) and HarmBenchHarmBench (right) prompt clustering for $5\text{-}CRI$, projected to $2D$ using t-SNE.
@@ B.4 CRI Integration Guidelines
To support the broad applicability of the $CRI$ framework, we outline generic integration strategies adaptable to various optimization-based paradigms. These strategies utilize $K\text{-}CRI$ templates to initialize prompts near the compliance subspace, enhancing efficiency while remaining agnostic to specific attack mechanisms. Integration approaches align and adaptable across gradient-based jailbreak attacks such sampling-based, and reinforcement-driven optimization settings. Concrete integration examples are provided in our experiments, in Section˜4.
@@
This paradigm employs gradient-based optimization to iteratively refine input prompts, minimizing the negative log probability of eliciting a target response. The $CRI$ framework integrates by initializing prompts utilizing $K$ pre-computed attacks, which serve as starting points for gradient descent, thereby reducing the number of iterations required for convergence. By leveraging the semantic clustering inherent in $CRI$, this approach ensures that initial prompts are optimally positioned within the compliance subspace, enhancing optimization efficiency Zou et al. (2023); Wallace et al. (2019); Shin et al. (2020); Guo et al. (2021); Ebrahimi et al. (2018). Similarly, $CRI$ can also be integrated for extended attack variants of this category, which utilize a scalar energy function Guo et al. (2024) to guide prompt modifications toward model outputs that align with desired behaviors, often without specifying a precise target response. $CRI$ integrates by selecting $K$ attacks that minimize the initial energy score, providing robust initial prompts for untargeted optimization. This strategy capitalizes on $CRI$’s ability to group prompts by semantic similarity, thereby improving generalizability across diverse prompt distributions Liu et al. (2023); Chao et al. (2023); Zhang et al. (2022); Wei et al. (2024); Li et al. (2023a). Results for Zou et al. (2023) can be found in Table˜1,
@@
Embedding attacks manipulate prompt representations within the latent space to circumvent refusal mechanisms, using training-time attacks as a baseline to identify vulnerable directions that trigger compliance. $CRI$ integrates by selecting $K$ attacks within the LLMs’ embedding space, followed by selecting the one that minimizes the attack criterion for a given prompt Leo et al. (2024); Qi et al. (2023); Li et al. (2024a); Zhao et al. (2023); Huang et al. (2023), results for Schwinn et al. (2024) appear in Table˜1.
@@
These attacks optimize a single transformation applicable across multiple prompts, prioritizing broad generalizability. $CRI$ integrates by training $K$ attacks on clustered subsets of the fine-tuning set, then selecting the optimal template for a given prompt during inference. This approach balances universal applicability with prompt-specific efficacy, leveraging $CRI$’s clustered structure to enhance transformation robustness Zou et al. (2023); Mozes et al. (2021); Liu et al. (2022); Brown et al. (2023); Yang et al. (2024b), results for Zou et al. (2023) can be found in Table˜3 for Llama-2 and Table˜3 for Vicuna.
@@
These attacks rely on querying the model without gradient access, using output feedback to iteratively refine prompts. $CRI$ integrates by employing $K$ attacks as an initial set, minimizing the number of model interactions required to achieve compliance. The diversity of $CRI$ clusters enhances query efficiency, particularly in resource-constrained black-box settings Zou et al. (2023); Chen et al. (2021); Garg et al. (2022); Li et al. (2023b); Zhang et al. (2024), for results on the GCGMGCG-M model transferability Zou et al. (2023) presented in Table˜3 for Llama-2-based jailbreak attack and Table˜3 for Vicuna-based jailbreak attack.
@@
Beyond standard optimization-based attacks, $CRI$ extends to additional paradigms such as prefix optimization, refusal unlearning, and alignment modulation Jones et al. (2024); Smith et al. (2024); Xu et al. (2023); Lin et al. (2023); Wu et al. (2024a). Its flexible $K$ initialization set framework enables adaptation to diverse objectives while remaining model- and method-agnostic. Moreover, $CRI$ integrates seamlessly with recent optimization improvements, including techniques like BOOST and I-GCG Boost et al. (2024); IGCG et al. (2024), enhancing convergence and success rates by leveraging robust initializations. An example integration of $CRI$ with BOOST, can be found in Figure˜27.
@@
Safety-Alignment divides input prompts into compliance and refusal regions—boundaries that adversaries can exploit (Marshall et al., 2024; Baumann, 2024; Huang et al., 2024; Yu et al., 2024c; Chao et al., 2023; Deng et al., 2023; Li et al., 2025; Yan et al., 2024; Xu et al., 2024b; Xiao et al., 2024a). Recent studies suggest refusal direction—a single vector in the residual stream that governs refusal behavior (Arditi et al., 2024). Modifying this direction can toggle the model’s safety responses, prompting offensive uses like CAST’s real-time content control (Lee et al., 2025) or defensive uses like GradSafe’s gradient-based detection (Xie et al., 2024). However, new work represents refusal cones—suggesting multiple interacting refusal directions (Wollschläger et al., 2025).
@@
Early Jailbreak attack approaches utilize uninformative seeds, employing repeated characters or random Gaussian token embeddings (Zou et al., 2023; Hu et al., 2024). Subsequent methods introduced techniques such as Langevin dynamics or beam search (Guo et al., 2024; Sadasivan et al., 2024; Mehrotra et al., 2023). Implicit initialization approaches rely on subtle mechanisms—few-shot distractors, puzzle-like games, or dynamic context shifts—that implicitly guide models toward generating harmful outputs (Xiao et al., 2024b; Chang et al., 2024; Ramesh et al., 2025). Transfer-based initializations leveraging previously successful jailbreak prompts as starting points for subsequent optimizations (Jia et al., 2024; Andriushchenko et al., 2024; Wu et al., 2025b). Explicit initialization methods directly embed crafted injection prompts or obfuscated content (Liu et al., 2024a, 2023; Schwinn et al., 2024; Liu et al., 2024b; Sun et al., 2025; Xu et al., 2025; Jiang et al., 2024; Zhao et al., 2025a). Additional possible enhancements to initialization strategies include employing energy-based objectives (Zhao et al., 2024), prefix-tuning methods (Zhu et al., 2024; Zhou et al., 2024a; Zhang and Wei, 2025; Li et al., 2024d), or gradient-based optimization techniques (Li et al., 2024b; Yu et al., 2024b).
@@ C.1 Experiments Settings
GCGGCG and the Embedding attacks keep the same settings as before.
@@ C.1.3 Attack Settings
For the $AutoDAN$ attack variants ($GA$, HGAHGA) Liu et al. (2023), we utilize the same reference file filled with hand-crafted jailbreaks, and perform $100$ optimization steps.